OpenStudy (rock_mit182):

• State the domain and range for each of your equations. Write them in interval notation. \[x = (y+2)^{2}\]

OpenStudy (rock_mit182):

@jim_thompson5910 @radar @SolomonZelman

OpenStudy (rock_mit182):

\[\sqrt{x}=\sqrt{(y+2)^{2} }\] \[\pm \sqrt{x} = y+2\]

OpenStudy (rock_mit182):

\[y = \sqrt{x}-2\]

OpenStudy (rock_mit182):